You are dealt one card out of a standard 52 card deck. What is the probability of being dealt a king or a numbered card?
Answer by stanbon(75887)

You are dealt one card out of a standard 52 card deck. What is the probability of being dealt a king or a numbered card?
------------------------
Numbered: 2,3,4,5,6,7,8,9
P(king or numbered) = P(king)+P(numbered)
------
= (4/52) + (32/52)
------
= 36/52
------
= 9/13
-------------
Cheers,
Stan H.